OpenStudy (anonymous):

estimate the angle between the two vectors <-2, 1> and <2,4>

geerky42 (geerky42):

Are you familiar with this formula? a⋅b = |a||b|cosθ

OpenStudy (anonymous):

no, can you please help

geerky42 (geerky42):

hmm, best way to estimate is to plot vector, where both vector start from same point, then guess angle from there.

geerky42 (geerky42):

Best approach I can think of...

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (anonymous):

what is the dot product and product of magnitudes?

OpenStudy (anonymous):

is that multiplying both x and adding then multiply both y and adding?

geerky42 (geerky42):

what class are you taking right now?

OpenStudy (anonymous):

math analysis

OpenStudy (jdoe0001):

hmmm

geerky42 (geerky42):

huh? is that after algebra 2?

OpenStudy (jdoe0001):

there's an assumption you'd know what a dot product and magnitude are

OpenStudy (anonymous):

yes, it it is the equivalent to pre calc

OpenStudy (anonymous):

I didn't learn vectors in algebra 2

geerky42 (geerky42):

ah i see, never heard of it being called math analysis haha.

OpenStudy (amistre64):

another route is trig <-2, 1> tan(a) = -1/2 and <2,4> tan(b) = 4/2 the difference between a and b is just: tan(a-b) = tan(a)-tan(b)/(1+tan(a)tan(b)) and invert the trig

OpenStudy (anonymous):

thank you, that makes more sense! I'm really trying to understand this, I appreciate your help and patience

OpenStudy (amistre64):

hmm, a zero divisor, what does that tell us? same thing as the dot product = 0 really

OpenStudy (anonymous):

perpendicular?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

which is the same as evaluating slopes: -1/2 *4/2 = -1 which is of course a property of perp slopes

OpenStudy (anonymous):

alright, thanks so much, I get it now!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thanks